OpenStudy (anonymous):

Hey i have a question about dividing radicals(square roots) I'm going to write the equation in the comments!

OpenStudy (anonymous):

\[5/(-5-3\sqrt{3})\] please show work if you can!

OpenStudy (jeyekomon):

I would try this: 5 / (-5-3*sqrt(3)) = -5 / (5+3*sqrt(3)) = -5 / (5+sqrt(27)) = (now a trick when we multiply it all by 1) = -5 / (5+sqrt(27)) * (5 - sqrt(27)) / (5 - sqrt(27)) = = (-5) * (5 - sqrt(27)) / (25 - 27) = 5 * (5 - sqrt(27)) / 2 Is that what you want? Hope it's clear how I did it..

OpenStudy (anonymous):

i just dont get how -5*5 is equal to -5.. shouldnt it be -25?

OpenStudy (jeyekomon):

And where exactly? Sorry, I can't see where I did 5*5..

OpenStudy (anonymous):

(-5) * (5 - sqrt(27)) / (25 - 27) = 5 * (5 - sqrt(27)) / 2 well i mean you did write -5*5 but could you simplify that to -25+3 sqrt3?

OpenStudy (anonymous):

or would it be -25-15sqrt3/2?

OpenStudy (jeyekomon):

(-5) * (5 - sqrt(27)) / (25 - 27) = = (-5) * (5 - sqrt(27)) / (-2) = = 5 * (5 - sqrt(27)) / 2 I made several steps together. It may be confusing..

OpenStudy (jeyekomon):

And yes, you can simplify that: 5 * (5 - sqrt(27)) / 2 = (25 - 15*sqrt(3)) / 2 but I don't think it's a big difference..

OpenStudy (anonymous):

ok thank you very much!